# Policy Optimization with Smooth Guidance Learned from State-Only Demonstrations

### v2

@@ 4.1 Smooth Guidance Reward Design
where $d(\cdot)=D(\cdot,\mathcal{M}_{E})$, and $\tau$ and $\upsilon$ have the same meanings as those in Eq. (9). $k$ is a predefined positive constant to adjust the value of $\omega$. Specifically, by choosing the suitable value of $k$, we can smooth the value of $\omega$ and prevent its value corresponding to a certain trajectory from being too large. It is worth noting that $D(\tau,\mathcal{M}_{E})$ is defined in Section 3.2. Similar to $p_{\pi}(\cdot)$, $\omega(\cdot|\mathcal{M}_{E})$ can be considered as a conditional probability density function over the trajectory space. $\omega(\cdot|\mathcal{M}_{E})$ satisfies the conditions of probability distribution $\omega_{\pi}(\tau|\mathcal{M}_{E})\geq 0,\forall\tau$ and $\int_{\tau}p_{\pi}(\tau)\omega(\tau|\mathcal{M}_{E})\mathrm{d}\tau=1$. According to $\omega_{\pi}(\tau|\mathcal{M}_{E})$, a trajectory closer to $\mathcal{M}_{E}$ is granted more weight by $\omega_{\pi}(\cdot|\mathcal{M}_{E})$. Based on the trajectory weight function $\omega_{\pi}(\tau|\mathcal{M}_{E})$, we can define the trajectory importance as follows:
@@
Here, $R(\cdot)$ denotes the return value of the trajectory. $\tau_{E}=\operatorname{arg\,min}_{\tau_{E}\in\mathcal{M}_{E}}\mathrm{MMD}^{2}(\tau,\tau_{E},\mathcal{H})$ is the state-only demonstration in $\mathcal{M}_{E}$ closest to $\tau$ under the MMD distance metric, and $\mathcal{H}$ a reproducing kernel Hilbert space as described in Section 3. The two constants $\alpha$ and $\beta$ in Eq. (12) are non-negative, and their linear sum is $1$, i.e., satisfies the condition $\alpha+\beta=1$. In this manner, the trajectory importance integrates the state distribution information of the demonstrations with the returns of the trajectories. Finally, we can define the smooth guidance rewards for a state-action pair $(s,a)$ as follows:
@@
Note that $x=f(s)$. The guidance reward $r_{i}(s,a)$ allocated to each state-action pair is the expectation of the product of $R_{j}(\tau)$ and $\omega(\tau|\mathcal{M}_{E})$ under the conditional distribution $p_{\pi}(\tau|x)$. The joint return $R_{j}(\tau)$ is the weighted average sum of the environmental returns of $\tau$ and $\tau_{E}$. Moreover, we use $\omega(\tau|\mathcal{M}_{E})$ to adjust its contribution to the guidance reward $r_{i}(s,a)$ according to Eq. (11). Specifically, the closer the trajectory is from $\mathcal{M}_{E}$, the larger the weight of $\tau$. Therefore, such a trajectory will contribute more to the smooth guidance reward. A new smoothed RL objective is obtained by inserting Eq. (9) in the standard RL objective:
@@
This is similar to the standard RL objective function in Eq. (1), albeit with the guidance reward different from the environmental rewards. By maximizing this objective, the proposed method encourages the agent to revisit the hopeful region of the state space. In this manner, the state-action distribution of the current policy will gradually align with that of the demonstration trajectories. Hence, The agent can finally generate highly rewarded trajectories similar to the demonstration trajectories in the replay memory. Furthermore, experimental results indicate that ascertaining the temporal structure of the sequential decision problem often results in high sample efficiency.
@@
4.2 Practical Algorithms
In practice, without no access to the true MDP dynamics, solving Eq. (13) for obtaining exact guidance rewards is infeasible. Fundamentally, $r_{i}(s,a)$ is the weighted expectation of joint rewards $R_{j}$ under the trajectory weight function $\omega(\tau|\mathcal{M}_{E})$. Hence, we resort to the Monte-Carlo (MC) method to calculate the estimation of the guidance reward: $r_{i}(s,a)=\mathbb{E}_{\tau\sim p_{\pi}(\tau|x)}\left[\omega(\tau|\mathcal{M}_{E})R_{j}(\tau)\right]$. Let $\mathcal{B}$ denote a trajectory buffer generated by the current policy $\pi$ in the MDP. Then the MC estimation of the guidance reward $r_{i}(s,a)$ can be given as:
@@ 4.2 Practical Algorithms
where $d(\cdot)=D_{\rm MMD}(\cdot,\mathcal{M}_{E})$ is introduced in Section 3, $k$ and $\epsilon$ are predefined positive constants.
@@
4.2.1 Applying to high-dimensional continuous spaces.
@@ 5 Theoretical Analysis
Suppose $\pi_{b}$ is a policy implied by the replay memory $\mathcal{M}_{E}$ which contains all optimal trajectories, and $p(s^{\prime}|s)=\pi_{b}(a|s)P(s^{\prime}|s,a)$ is the state transition function consistent with $\mathcal{M}_{E}$. Let the discount factor $\lambda$ be 1. According to the definition of the smooth reward $r_{i}(s,a)$, if the current policy $\pi$ is expressed as:
[…]
@@
Next, a deeper look at the proposed POSG algorithm is given to demonstrate its effectiveness in improving control performance. We quantitatively analyze the performance improvement bound in each iteration. The result obtained in this study indicates that POSG has a solid performance improvement guarantee compared to ordinary DRL algorithms. This conclusion relates the theoretical analysis of POSG with the results in previous works kakade2002approximately ; Schulman2015TrustRP ; achiam2017constrained .
@@
Before proceeding, it is worth mentioning that Theorem 1 is similar to Corollary 1 of  achiam2017constrained . In the above theorem, a new performance improvement bound $(1+\lambda)D_{\pi,f}(\pi^{\prime})$ is derived based on the environmental advantage function $A_{e}(s,a)$ and the smooth guidance reward function $r_{i}(s,a)$. This result illustrates that the smooth guidance reward can result in a broader performance improvement range with higher upper and lower bounds than that obtained only with the environmental reward. Consequently, the proposed method allows the agent to obtain a larger performance improvement in each policy optimization iteration.
Note that the bound in Theorem 1 we have given is based on TV divergence between policies. For ease of calculation, we further connect the performance bound to KL divergence through Pinsker’s inequality csiszar2011information : for two arbitrary distributions $p,q$, the TV and KL divergences satisfy the following in-equation: $D_{\rm TV}(p\parallel q)\leq\sqrt{D_{\rm KL}(p\parallel q)/2}$. Combining this with Jensen’s in-equation, we obtain
$\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm TV}(\pi^{\prime}\parallel\pi)[s]\right]\leq\underset{s\sim d^{\pi}}{\mathbb{E}}\left[\sqrt{\frac{1}{2}D_{\rm KL}(\pi^{\prime}\parallel\pi)[s]}\right]\leq\sqrt{\frac{1}{2}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm KL}(\pi^{\prime}\parallel\pi)[s]\right]}.$ (20)
@@
The proposed guidance reward can be viewed as a smooth credit assignment over the trajectory space and characterize the influence and impact of each state-action pair the agent performs on future rewards. When the smooth guidance reward is used to update the current policy, the worst-case lower bound described in this corollary is $1+\lambda$ times the previous result obtained by achiam2017constrained . This result indicates that the proposed method generates a higher performance improvement boundary and provides a more stable performance guarantee in each iteration. Thus, Corollary 1 theoretically demonstrates the crucial advantage of POSG in obtaining higher control performances and accelerating training.
6 Experimental Setup
6.1 Environments
@@
Fig. 1: A collection of environments that we used to evaluate POSG: (a) Key-Door-Treasure domain; (b) SparseHalfCheetah; (c) SparseHopper; (d) Ant Maze.
@@
Hard-Exploration Ant Maze. To further investigate the performance of POSG, we evaluated this algorithm with the ant maze task introduced as the benchmark for RL by duan2016benchmarking , which is depicted in Fig. 1d. The hard-exploration property of this task is caused by two factors: continuous locomotion control and navigation in the maze. More specifically, the Ant robot must first learn to walk smoothly and then struggle to reach the target position in the maze. The Ant robot is only rewarded a large positive bonus when it reaches the specified position of the maze. Meanwhile, the episode will end when the agent receives the reward. The state space of this environment consists of two main parts: the agent’s internal joint angle information and task-specific attributes. The agent’s joint angle information is only determined by the agent’s internal state, and the task-specific attributes are mainly obtained from sensor readings, including the positions of walls and goals.
6.2 Neural Architectures and Hyper-parameters
@@
6.3 Baseline Methods
To investigate the benefits of using the smooth guidance rewards in the sparse reward setting, several different baseline methods were used for performance comparison in different tasks. We first compared our method with the state-of-the-art RL methods that learn from state-only demonstrations. To achieve this goal, we adopted GAIfO torabi2018generative as the baseline. However, this method is a state-only imitation learning method, and it does not interact with the environment to gain new experiences during training. For a fair comparison, we combined GAIfO with SIL oh2018self and obtained a new RL method GASIfO. Furthermore, we used several other state-of-the-art RL baselines, including SIL, PPO+D libardi2021guided and generative adversarial self-imitation learning (GASIL) guo2018generative . Unlike the standard SIL and GASIL framework in guo2018generative , our implemented GASIL could obtain the same demonstrations as POSG at the beginning of training. It is worth that these three methods can access the demonstrator’s action information. We also ran proximal policy optimization (PPO) schulman2017proximal baseline to verify the reward sparsity. For all baseline approaches, we adopted the parameters that produced the best performance during the parameter search. All performance curves were obtained by averaging over ten separate runs with different random seeds, and the shaded regions represented the standard error over these ten runs.
6.4 Acquisition of offline state-only demonstrations
@@
7 Evaluation of Results
In this section, to thoroughly assess the proposed POSG algorithm, we test it in three different classes of environments introduced in Section 6, respectively, and compare the performance with other baseline algorithms. The experimental results demonstrate that POSG can achieve better performance than other methods.
7.1 Evaluations in the Key-Door-Treasure domain
7.1.1 Performance Comparison with Baseline Methods
@@
Fig. 2: (a) Success rate in the Key-Door-Treasure domain; (b) The changing trend of the MMD distance; (c) State-action visitation graph of demonstrations; (d) State-action visitation graph of the POSG learned policy.
We further investigated the changing trend of the MMD distance for all methods during the training process. The changing curves of the MMD distance are depicted in Fig. 2b. As shown in Fig. 2b, the value of the MMD distance decreases to zero at the end of the training for the POSG, SIL, and PPO+D algorithms. This result indicates that these three approaches acquire the optimal behaviors that a demonstrator recommends. Furthermore, POSG’s MMD distance reduction rate is the highest among all methods. This result suggests that POSG avoids excessive meaningless exploration, and the smooth guidance reward provides more efficient policy optimization directions. Meanwhile, we draw the state-action visitation graphs of demonstrations and POSG learned policy in Figs. 2c and 2d, respectively. The comparative result demonstrated that the POSG agent imitates experts’ behavior almost perfectly.
7.1.2 Visualization Comparisons for Exploration
@@
Fig. 3: The state-action visitation graphs of all algorithms: (a) PPO; (b) SIL; (c) PPO+D; (d) POSG.
@@
7.2 Performance on MuJoCo Locomotion Tasks
7.2.1 Performance Comparisons with Baseline Methods
@@
Figs. 4a and 4b show that POSG outperforms other baseline approaches in the locomotion control tasks. Specifically, POSG can converge faster than different algorithms during policy optimization and achieve a higher final return. In contrast, PPO+D policy gradient computation only relies on the advantage information of demonstrations and thus cannot perform effective policy optimization with sparse-reward demonstrations. This causes the performance difference between PPO+D and POSG, and even the average return of PPO+D on the SparseHopper task no longer increases from the beginning of training. Since GASIL and GASIfO ignore reward signals of demonstrations and solely care about the distribution difference between the current policy and the sparse-reward demonstration data, they cannot obtain competitive results compared to POSG. The performance difference between GASIL and GASIfO in Fig. 4a demonstrates the importance of action information for learning. These results show that POSG achieves simple and efficient credit assignments with sparse-reward demonstrations. POSG avoids information bottlenecks by integrating the sparse-reward demonstrations’ distribution information with the relevant trajectories’ returns.
@@
Fig. 4: (a) Learning curves of average return on the SparseHalfCheetah task; (b) Learning curves of average return on the SparseHopper task; (c) Learning curves of average ground truth reward per step on the SparseHopper task. The dotted line represents the return value of the demonstration trajectories under the ground truth reward function.
7.2.2 Evaluation of Ground-Truth Reward Learning

### v3

@@ IV-B Trajectory Importance Evaluation Mechanism
where $d(\cdot)=D(\cdot,\mathcal{M}_{E})$, $\tau$ and $\upsilon$ have the same meanings as those in Eq. (7), and $p_{\pi}(\cdot)$ is the trajectory distribution defined in Eq. (6). $k$ is a predefined positive constant to adjust the value of $\omega$. Specifically, by choosing the suitable value of $k$, we can smooth the value of $\omega$ and prevent its value corresponding to a certain trajectory from being too large or too small. Similar to $p_{\pi}(\cdot)$, $\omega(\cdot|\mathcal{M}_{E})$ can be considered as a conditional probability density function over the trajectory space. $\omega(\cdot|\mathcal{M}_{E})$ satisfies the conditions of probability distribution $\omega_{\pi}(\tau|\mathcal{M}_{E})\geq 0,\forall\tau$ and $\int_{\tau}p_{\pi}(\tau)\omega(\tau|\mathcal{M}_{E})\mathrm{d}\tau=1$. According to $\omega_{\pi}(\tau|\mathcal{M}_{E})$, a trajectory closer to $\mathcal{M}_{E}$ is granted more weight by $\omega_{\pi}(\cdot|\mathcal{M}_{E})$. Based on the trajectory weight function $\omega_{\pi}(\tau|\mathcal{M}_{E})$, we can define the trajectory importance as follows:
@@
Here, $R(\cdot)$ denotes the return value of the trajectory. $\tau_{E}=\operatorname{arg\,min}_{\tau_{E}\in\mathcal{M}_{E}}\mathrm{MMD}^{2}(\tau,\tau_{E},\mathcal{H})$ is the state-only demonstration in $\mathcal{M}_{E}$ closest to $\tau$ under the MMD distance metric, and $\mathcal{H}$ a reproducing kernel Hilbert space as described in Section III. The two constants $\alpha$ and $\beta$ in Eq. (13) are non-negative, and their linear sum is $1$, i.e., satisfies the condition $\alpha+\beta=1$. In this manner, the importance of the trajectory integrates the state distribution information of the demonstrations with the returns of the trajectories.
IV-C Smooth Guidance Reward Computation Technology
In this section, we define the smooth guidance rewards for a state-action pair $(s,a)$ as follows:
@@ IV-C Smooth Guidance Reward Computation Technology
Here, $p_{\pi}(\cdot|x)$ is the probability density model for trajectories including $x$ defined in Eq. (7). Note that $x=f(s)$. The guidance reward $r_{i}(s,a)$ allocated to each state-action pair is the expectation of the product of $R_{j}(\tau)$ and $\omega(\tau|\mathcal{M}_{E})$ under the conditional distribution $p_{\pi}(\tau|x)$. The joint return $R_{j}(\tau)$ is the weighted average sum of the environmental returns of $\tau$ and $\tau_{E}$. Moreover, we use $\omega(\tau|\mathcal{M}_{E})$ to adjust its contribution to the guidance reward $r_{i}(s,a)$ according to Eq. (12). Specifically, the closer the trajectory is from $\mathcal{M}_{E}$, the larger the weight of $\tau$. Therefore, such a trajectory will contribute more to the smooth guidance reward. A new smoothed RL objective is obtained by inserting Eq. (14) in the standard RL objective:
@@
This is similar to the standard RL objective function in Eq. (1), albeit with the guidance reward different from the environmental rewards. By maximizing this objective, the proposed method encourages the agent to revisit the hopeful region of the state space. In this manner, the state-action distribution of the current policy will gradually align with that of the demonstration trajectories. Hence, the agent can finally generate highly rewarded trajectories similar to the demonstration trajectories in the replay memory. Furthermore, experimental results indicate that determining the temporal structure of the sequential decision problem often results in high sample efficiency.
@@
IV-D Algorithm Implementations for Discrete and Continuous Control Tasks
In practice, without access to the true MDP dynamics, solving Eq. (14) to obtain exact guidance rewards is infeasible. Fundamentally, $r_{i}(s,a)$ is the weighted expectation of joint rewards $R_{j}$ under the trajectory weight function $\omega(\tau|\mathcal{M}_{E})$. Hence, we resort to the Monte Carlo (MC) method to calculate the estimation of the guidance reward: $r_{i}(s,a)=\mathbb{E}_{\tau\sim p_{\pi}(\tau|x)}\left[\omega(\tau|\mathcal{M}_{E})R_{j}(\tau)\right]$. Let $\mathcal{B}$ denote a trajectory buffer generated by the current policy $\pi$ in the MDP. Then, the MC estimation of the guidance reward $r_{i}(s,a)$ can be given as:
@@ IV-D Algorithm Implementations for Discrete and Continuous Control Tasks
where $d(\cdot)=D_{\rm MMD}(\cdot,\mathcal{M}_{E})$ is introduced in Section III, $k$ and $\epsilon$ are predefined positive constants.
@@
IV-D1 Applying to high-dimensional continuous spaces.
@@ V Theoretical Analysis
Suppose $\pi_{b}$ is a policy implied by the replay memory $\mathcal{M}_{E}$ that contains all optimal trajectories, and $p(s^{\prime}|s)=\pi_{b}(a|s)P(s^{\prime}|s,a)$ is the state transition function consistent with $\mathcal{M}_{E}$. Let the discount factor $\lambda$ be 1. According to the definition of the smooth reward $r_{i}(s,a)$, if the current policy $\pi$ is expressed as:
@@
Next, a deeper look at the proposed POSG algorithm is given to demonstrate its effectiveness in improving control performance. We quantitatively analyze the performance improvement bound in each iteration. The result obtained in this study indicates that POSG has a solid performance improvement guarantee compared to ordinary DRL algorithms. This conclusion relates the theoretical analysis of POSG with the results in previous works [61, 6, 62].
@@
Figure 1: A collection of environments that we used to evaluate POSG: (a) Key-Door-Treasure domain; (b) SparseHalfCheetah; (c) SparseHopper; (d) Ant Maze.
@@
Before proceeding, it is worth mentioning that Theorem 1 is similar to Corollary 1 of  [62]. In the above theorem, a new performance improvement bound $(1+\lambda)D_{\pi,f}(\pi^{\prime})$ is derived based on the environmental advantage function $A_{e}(s,a)$ and the smooth guidance reward function $r_{i}(s,a)$. This result illustrates that the smooth guidance reward can result in a broader performance improvement range with higher upper and lower bounds than that obtained only with the environmental reward. Consequently, the proposed method allows the agent to obtain a larger performance improvement in each policy optimization iteration.
Note that the bound in Theorem 1 we have given is based on TV divergence between policies. For ease of calculation, we further connect the performance bound to KL divergence through Pinsker’s inequality [63]: for two arbitrary distributions $p,q$, the TV and KL divergences satisfy the following in-equation: $D_{\rm TV}(p\parallel q)\leq\sqrt{D_{\rm KL}(p\parallel q)/2}$. Combining this with Jensen’s in-equation, we obtain
$\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm TV}(\pi^{\prime}\parallel\pi)[s]\right]$ $\leq\underset{s\sim d^{\pi}}{\mathbb{E}}\left[\sqrt{\frac{1}{2}D_{\rm KL}(\pi^{\prime}\parallel\pi)[s]}\right]$ (21)
$\leq\sqrt{\frac{1}{2}\underset{s\sim d^{\pi}}{\mathbb{E}}\left[D_{\rm KL}(\pi^{\prime}\parallel\pi)[s]\right]}.$
@@
The proposed guidance reward can be viewed as a smooth credit assignment over the trajectory space and characterize the influence and impact of each state-action pair the agent performs on future rewards. When the smooth guidance reward is used to update the current policy, the worst-case lower bound described in this corollary is $1+\lambda$ times the previous result obtained by [62]. This result indicates that the proposed method generates a higher performance improvement boundary and provides a more stable performance guarantee in each iteration. Thus, Corollary 1 theoretically demonstrates the crucial advantage of POSG in obtaining higher control performances and accelerating training.
VI Experimental Setup
VI-A Environments
@@
Hard-Exploration Ant Maze. To further investigate the performance of POSG, we evaluated this algorithm with the ant maze task introduced as the benchmark for RL by [64], which is depicted in Fig. 1d. The hard-exploration property of this task is caused by two factors: continuous locomotion control and navigation in the maze. More specifically, the Ant robot must first learn to walk smoothly and then struggle to reach the target position in the maze. The Ant robot is only rewarded a large positive bonus when it reaches the specified position of the maze. Meanwhile, the episode will end when the agent receives the reward. The state space of this environment consists of two main parts: the agent’s internal joint angle information and task-specific attributes. The agent’s joint angle information is only determined by the agent’s internal state, and the task-specific attributes are mainly obtained from sensor readings, including the positions of walls and goals.
VI-B Neural Architectures and Hyper-parameters
@@
VI-C Baseline Methods
To investigate the benefits of using the smooth guidance rewards in the sparse reward setting, several different baseline methods were used for performance comparison in different tasks. We first compared our method with the state-of-the-art RL methods that learn from state-only demonstrations. To achieve this goal, we adopted GAIfO [43] as the baseline. However, this method is a state-only imitation learning method, and it does not interact with the environment to gain new experiences during training. For a fair comparison, we combined GAIfO with SIL [33] and obtained a new RL method GASIfO. Furthermore, we used several other state-of-the-art RL baselines, including SIL, PPO+D [27] and generative adversarial self-imitation learning (GASIL) [65]. Unlike the standard SIL and GASIL framework in [65], our implemented GASIL could obtain the same demonstrations as POSG at the beginning of training. It is worth that these three methods can access the demonstrator’s action information. We also ran proximal policy optimization (PPO) [42] baseline to verify the reward sparsity. For all baseline approaches, we adopted the parameters that produced the best performance during the parameter search. All performance curves were obtained by averaging over ten separate runs with different random seeds, and the shaded regions represented the standard error over these ten runs.
VI-D Acquisition of offline state-only demonstrations
@@
Figure 2: (a) Success rate in the Key-Door-Treasure domain; (b) The changing trend of the MMD distance.
VII Evaluation of Results
In this section, to thoroughly assess the proposed POSG algorithm, we test it in three different classes of environments introduced in Section VI, respectively, and compare the performance with other baseline algorithms. The experimental results demonstrate that POSG can achieve better performance than other methods.
Figure 3: (a) State-action visitation graph of demonstrations; (b) State-action visitation graph of the POSG learned policy.
VII-A Evaluations in the Key-Door-Treasure domain
VII-A1 Performance Comparison with Baseline Methods
@@
We further investigated the changing trend of the MMD distance for all methods during the training process. The changing curves of the MMD distance are depicted in Fig. 2b. As shown in Fig. 2b, the value of the MMD distance decreases to zero at the end of the training for the POSG, SIL, and PPO+D algorithms. This result indicates that these three approaches acquire the optimal behaviors that a demonstrator recommends. Furthermore, POSG’s MMD distance reduction rate is the highest among all methods. This result suggests that POSG avoids excessive meaningless exploration, and the smooth guidance reward provides more efficient policy optimization directions. Meanwhile, we draw the state-action visitation graphs of demonstrations and POSG learned policy in Figs. 3a and 3b, respectively. The comparative result demonstrated that the POSG agent imitates experts’ behavior almost perfectly.
VII-A2 Visualization Comparisons for Exploration
@@
Figure 4: The state-action visitation graphs of all algorithms: (a) PPO; (b) SIL; (c) PPO+D; (d) POSG.
@@
VII-B Performance on MuJoCo Locomotion Tasks
VII-B1 Performance Comparisons with Baseline Methods
@@
Figure 5: (a) Learning curves of average return on the SparseHalfCheetah task; (b) Learning curves of average return on the SparseHopper task.
@@
Figs. 5a and 5b show that POSG outperforms other baseline approaches in the locomotion control tasks. Specifically, POSG can converge faster than different algorithms during policy optimization and achieve a higher final return. In contrast, PPO+D policy gradient computation only relies on the advantage information of demonstrations and thus cannot perform effective policy optimization with sparse-reward demonstrations. This causes the performance difference between PPO+D and POSG, and even the average return of PPO+D on the SparseHopper task no longer increases from the beginning of training. Since GASIL and GASIfO ignore reward signals of demonstrations and solely care about the distribution difference between the current policy and the sparse-reward demonstration data, they cannot obtain competitive results compared to POSG. The performance difference between GASIL and GASIfO in Fig. 5a demonstrates the importance of action information for learning. These results show that POSG achieves simple and efficient credit assignments with sparse-reward demonstrations. POSG avoids information bottlenecks by integrating the sparse-reward demonstrations’ distribution information with the relevant trajectories’ returns.
@@
Figure 6: Learning curves of average ground truth reward per step on the SparseHopper task. The dotted line represents the return value of the demonstration trajectories under the ground truth reward function.
VII-B2 Evaluation of Ground-Truth Reward Learning